If m<2 = x/2 -10 and m<3= x/3 + 40, find x and m<2
Note: these angles are supplementary

Answer by stanbon(75887)

If m<2 = x/2 -10 and m<3= x/3 + 40, find x and m<2
Note: these angles are supplementary
-----
Eq::
(x/2-10) + (x/3+40) = 180 degrees
-----
(5/6)x + 30 = 180
----
x = 6/5 * 150
-----
x = 180
----
m<2 =(x/2) - 10 = 180/2 - 10 = 80 degrees
---------------
Cheers,
Stan H.
--------------